# Using Large Language Models for Hyperparameter Optimization

### v1

@@ 4 Results
We report our results for tuning hyperparameters on eight different datasets and four models (neural nets, SVMs, random forests, and logistic regression) from the HPOBench benchmark for hyperparameter optimization (Eggensperger et al., 2021). We then analyze low-dimensional optimization problems to understand and visualize the LLM trajectories in a simpler setting and then discuss how our approach can be applied to code generation.
@@ 4.1 Hyperparameter Optimization on HPOBench
We report results on the HPOBench (Eggensperger et al., 2021) benchmark, which defines search spaces for various models on publicly available OpenML AutoML benchmarks. Eggensperger et al. (2021) show that the benchmarks yield different hyperparameter-loss landscapes and are effective for benchmarking algorithm performance. We used the first $8$ datasets provided and tuned hyperparameters for the algorithms implemented in Sklearn: logistic regression, support vector machines (SVMs), random forests, and neural networks (Pedregosa et al., 2011), resulting in $32$ different tasks. We tune $2$ hyperparameters for SVMs, $2$ for logistic regression, $4$ for random forests, and $5$ for neural networks (Eggensperger et al., 2021). Search space details can be found in Appendix B, as well as figures of results on each task. We compare the following HPO approaches:
@@
LLMs. We use the most recent time-stamped versions of GPT-3.5 and GPT-4 models from OpenAI in our initial experiments: gpt-3.5-turbo-0613 and gpt-4-0613 with temperature 0. Since GPT-4 Turbo was released after our initial experiments, we also conducted experiments with the model gpt-4-1106-preview and found it performed better at less than half the price of GPT-4.222OpenAI says “preview” means the gpt-4-1106-preview model is not suited for production traffic. We did not encounter any problems in our experiments.
@@
Random. Random search often outperforms grid and manual search, thriving when the hyperparameter space has low effective dimensionality (Bergstra and Bengio, 2012). The design spaces of a given benchmark dictate where we should sample – e.g., log-uniformly for regularization hyperparameters in $\mathbb{R}^{+}$ or from a discrete set such as the number of neural network layers. We randomly sample $500$ configurations for each model and dataset combination, evaluate the corresponding losses, and then use a bootstrapped sample to estimate the loss on a given budget.
@@
Bayesian Optimization (BO). We evaluate two Bayesian HPO methods in the SMAC3 library (Lindauer et al., 2022) that use random forest and Gaussian processes as surrogate models.
In Table 1, we report aggregated results comparing two LLMs, random search, and two BO algorithms with different surrogate models. We compare the difference in validation error versus random search ($0.2$ to $0.1$ is a $50\%$ improvement). Every algorithm has a search budget of $10$ function evaluations. As a summary metric, we report the mean and median improvement in the $32$ tasks of each algorithm versus random search. Full results detailing the performance of each dataset and model are in Figure 2 and Figure 6. GPT-4 Turbo beats random search most frequently and achieves the highest mean and median improvement. When we compare the mean rank for the five search algorithms across the $32$ tasks – again, GPT-4 Turbo is consistently better.
@@
Table 1: We summarize the performance of various HPO Algorithms versus random search on $8$ datasets and $4$ search spaces tuning hyperparameters for logistic regression, SVMs, random forests, and neural networks. Bayesian optimization uses Gaussian process ($GP$) and random forest ($RF$) surrogate models. For all tasks, we use $10$ function evaluations. As summary metrics, we report how often each method beats random search (GPT-4 Turbo beats random search 81.25% of the time). We also compute the change in validation error for each optimizer versus random search and report the median and mean change across $32$ tasks. The mean rank is computed between the $5$ HPO approaches and random search, i.e. each method is assigned a rank between $1$ and $6$ on a task. The mean rank for random across the 32 tasks is $4.00$. Figure 6 shows the detailed performance on the other tasks.
@@
Model Beats Random ($\uparrow$) Median Change ($\uparrow$) Mean Change ($\uparrow$) Mean Rank ($\downarrow$)
@@
Table 2: Chain-of-thought reasoning prompting may offer modest improvement versus prompting a model to output the next set of hyperparameters immediately.
@@
Model Beats Random ($\uparrow$) Median Change ($\uparrow$) Mean Change ($\uparrow$) Rank ($\downarrow$)
@@
In Table 2, we evaluate the effects of chain-of-thought (Wei et al., 2022) on performance. To elicit reasoning, our intermediate messages to the language model are:
@@
The loss is expressed in scientific notation to four decimal places. See Appendix B for complete details on the prompts. We see that including reasoning has a large positive impact on GPT-3.5 and a marginal impact on GPT-4. Reasoning chains can be informative for a practitioner at the cost of additional tokens. Below is a snippet of GPT-4 reasoning when tuning hyperparameters for a neural network.
  • (loss $0.193$) GPT-4: The loss remained the same despite changes in the hyperparameters. This suggests that the model might not be sensitive to these parameters or has reached a local minimum. We should try a more drastic change in the learning rate and depth. <proposed config consistent with reasoning>
@@
We provide additional full trajectories in the Appendix F. We qualitatively observe reasoning that trades off between exploration and selecting a strong final value.
@@
We additionally evaluate our approach on a longer time horizon, with $30$ function queries, and show the results in Table 3. Here, we use the “compressed” prompt described in Section 2: we prompt the LLM with a single message at each step, showing the entire history of hyperparameter configurations and losses. A longer search trajectory has a reduced dependency on initialization and requires a better search algorithm for good performance. GPT-4 Turbo still compares favorably with other approaches. Qualitatively, the reasoning generated by the language model is a useful summary of experimental results and the relative importance of different hyperparameters. For example, when tuning a random forest, we get the following reasoning:
@@
Additionally, we run a preliminary experiment to assess if LLMs can be useful to initialize Bayesian optimization – we find that for a search trajectory of length 30, using the LLM proposed configurations for the first ten steps improves or matches performance on 21 of the 32 tasks ($65.6\%$) for Bayesian optimization with a random forest surrogate function.
@@ 4.2 Code generation
Code generation offers a flexible paradigm to specify the configuration space for hyperparameter tuning. In this setup, we treat the model and optimizer source code as hyperparameters tuned by the LLM. We use code generation as a method to circumvent requiring a specific configuration space to search over. The code generation approach is evaluated against random search and LLMs with a fixed, configuration-based search space. The baseline methods are provided with a space, which includes the depth and width of the network, the batch size, learning rate, and weight decay. To minimize the risk of data leakage issues where the LLM was trained on good hyperparameter settings for our dataset, we used a recent Kaggle NYC Taxi dataset Taxi and Commission (2023a, b) for our experiments. We provide our prompts and additional details in Appendix C.
On the first tuning iteration, we ask the LLM (gpt-4-1106-preview) to output source code for the model and optimizer, which implicitly also requires specifying the hyperparameters. This code is generated as functions that take in hyperparameters as arguments. In later iterations, the hyperparameter tuning is done by asking the LLM to generate functions calls with specific arguments. The model is run with the given settings, and the LLM receives the validation loss and the average training losses during each epoch as feedback.
Table 4 reports the minimum test loss achieved using a fixed search budget of $5$ evaluations, which simulates the results of the initial search phase during hyperparameter tuning. We randomly sample 200 configurations and use a bootstrapped sample to estimate the standard error for a random search. Figure 4 reports the minimum test loss found at each step and illustrates that code generation obtains better initial settings than competing approaches.
Figure 4: Test loss trajectory comparison with a standard error displayed over $5$ runs. We randomly sample 200 configurations and use a bootstrapped sample to estimate the standard error for a random search. Code generation and configuration-based search both compare favorably to random search.
@@
4.3 2-Dimensional Landscapes
To evaluate whether LLMs can reason about optimization choices, we study a set of 2-dimensional toy test functions commonly used in optimization: Rosenbrock (Rosenbrock, 1960), Branin (Dixon, 1978), Himmelblau (Himmelblau et al., 2018), and Ackley (Ackley, 1987). For each test function $f$, we consider optimizing $f(\boldsymbol{x})$ and also $f(\boldsymbol{x}-\boldsymbol{c})$, where $\boldsymbol{c}$ is a fixed constant, with $\boldsymbol{c}_{i}\sim\mathcal{U}(0,1)$, to mitigate the LLM’s potential to memorize the original $f$. Also, we benchmark performance on a well-conditioned and ill-conditioned quadratic. We provide the experimental details and the explored prompts in Appendix D.
We search for optimal solutions using LLMs as an optimizer with GPT-4. While these problems can be considered black-box optimization, we can also pose them as HPO to the LLM, using the same prompts in Figure 1 for consistency. We show the performance of a single fixed seed and temperature 00 in Figure 5. Performance across $3$ seeds outperforms random search in most tasks, as shown in Table 7. Moreover, inspecting the trajectories and reasoning chains shown in Appendix F, we qualitatively observe different behaviors that may characterize a strong search algorithm: initial exploration around the center and boundaries of the search space; a line-search type algorithm to probe function behavior; and reasoning that trades off between exploration and selecting a performant final value. However, performance can be inconsistent between random seeds, as shown in Table 6.
We use prompts with minimal details to gain intuition on this HPO approach and to evaluate our LLM approach on problems with less complexity. Inspecting the trajectories, we see room for more specific prompts. For example, the LLMs re-explored previously evaluated points. We obtain the same or better performance in $8$ of the $10$ tasks by adding a sentence stating that the functions are deterministic, shown in Table 8. We speculate that HPO performance can generally be improved with more details on the problem setting in the prompt.
Figure 5: The LLM optimizer effectively optimizes the function in most situations with few function evaluations, even when we apply random argument shifts to mitigate memorization.
@@ 5 Discussion
When performing HPO, LLMs can extend beyond fixed configuration spaces for hyperparameter tuning, as they can offer interactive help to debug and improve models using natural language. We highlight some limitations and potential applications.
5.1 Limitations
Limited Context Length.
A primary limitation of our prompting-based approach is the limited context length. LLMs typically have a predetermined context window size and cannot process prompts exceeding this limit in the token count. As we incorporate the history of previously tried suggestions from the LLMs, the number of input tokens in the prompt increases with each additional round of hyperparameter optimization and the total number of hyperparameters being tuned. Allocating more budgets to the hyperparameters requires longer context lengths. Advancements in research aimed at extending the context length for LLMs (Chen et al., 2023b) can potentially improve results by allowing a higher budget for hyperparameter optimization. Indeed, we see that GPT-4-Turbo already has a context length of 128k, significantly longer than previous OpenAI models.
@@ 5.1 Limitations
The exact inference procedures for LLMs like GPT-4 (OpenAI, 2023) are not publicly disclosed, making replicating the results in this work potentially challenging. For example, although GPT-4 should ideally give the same results when the sampling temperature is set to 00, this is not always true. For future studies, using open source LLMs (Taori et al., 2023; Chiang et al., 2023; Touvron et al., 2023a, b) to establish a benchmark with reproducible results would be beneficial. We focus on the current most general-purpose capable models, as most open-source models perform worse than GPT-3.5 on benchmarks.
Possibility of Dataset Contamination.
It is possible that some of the evaluation benchmarks used were part of the LLM training data. However, previous research suggests that such contamination might not critically influence overall performance (Radford et al., 2021). Building more private data sets unseen by LLM during training would be valuable for future studies or even synthetic datasets like our quadratic experiments. Additionally, we emphasize that our prompting process only provides the output of a function and does not contain any information about the benchmark, reducing the likelihood of exact memorization.
@@
It can be costly to conduct a full-scale HPO experiment with GPT-4. For example, executing $32$ tasks using chain-of-thought reasoning on HPOBench costs approximately \$$\$8$ USD as of September 2023. This cost can increase further when using more budget, but can be reduced by using the “compressed" prompt. We believe this burden will be alleviated as the models improve to be both capable and cost-efficient.
5.2 LLMs as a Hyperparameter Tuning Assistant
We run initial experiments to demonstrate the potential of an LLM as a general-purpose assistant for training pipelines. In Figure 7, LLMs provide useful feedback for the error messages, which is infeasible with traditional approaches. However, this can suffer from the challenges that affect current language models, such as hallucinations (Bowman, 2023). In Figure 8, the LLM assumes that the user has a neural network and suggests regularization, which may be unhelpful if the model is underfitting. Exploring this paradigm further is an exciting direction for future work.

### v2

@@ 4 Results
We evaluated our approach through a series of experiments. First, we assessed performance using the HPOBench benchmark (Eggensperger et al., 2021), where we tune hyperparameters across eight datasets for four different models: neural networks, SVMs, random forests, and logistic regression. We then used LLMs to tune hyperparameters for more complex architectures by optimizing Vision Transformers and ResNets on CIFAR-10. We also demonstrate how our method generalizes to code generation tasks. To build intuition about how LLMs navigate the optimization landscape, we analyzed their behavior on CIFAR-10 and low-dimensional problems in Appendix E, where we can directly visualize the optimization trajectories.
@@ 4.1 Hyperparameter Optimization on HPOBench
We report results on the HPOBench (Eggensperger et al., 2021) benchmark, which defines search spaces for various models on publicly available OpenML AutoML benchmarks. Eggensperger et al. (2021) showed that their benchmarks yield different hyperparameter-loss landscapes and are effective for benchmarking algorithm performance. We used the first $8$ datasets provided and tuned hyperparameters for the algorithms implemented in Sklearn: logistic regression, support vector machines (SVMs), random forests, and neural networks (Pedregosa et al., 2011), resulting in $32$ different tasks. As defined by the benchmark, we tune $2$ hyperparameters for SVMs, $2$ for logistic regression, $4$ for random forests, and $5$ for neural networks. Search space details can be found in Appendix B, as well as figures of results on each task. We compare the following HPO approaches:
@@
LLMs. Our initial experiments use OpenAI’s time-stamped versions of GPT-3.5 and GPT-4 models in: gpt-3.5-turbo-0613 and gpt-4-0613 with temperature 00. GPT-4 Turbo was released after our initial results, so we conducted experiments with the model gpt-4-1106-preview and found better performance at less than half the price of GPT-4.222OpenAI says “preview” means gpt-4-1106-preview is unsuited for production traffic. We did not encounter problems.
@@
Random. Random search often outperforms grid and manual search, thriving when the hyperparameter space has low effective dimensionality (Bergstra & Bengio, 2012). The design spaces of a given benchmark dictate where we should sample – e.g., log-uniformly for regularization hyperparameters in $\mathbb{R}^{+}$ or from a discrete set such as the number of neural network layers.
@@
Bayesian Optimization (BO). We evaluate two Bayesian HPO methods in the SMAC3 library (Lindauer et al., 2022) that use random forest and Gaussian processes as surrogate models. We use the default recommended settings, including an initial design of up to $25\%$ of the optimization budget.
@@
Table 1: We summarize the performance of various HPO Algorithms on $8$ datasets and $4$ search spaces tuning hyperparameters for logistic regression, SVMs, random forests, and neural networks. For all tasks, we use $10$ function evaluations. As summary metrics, we report how often each method beats random search (GPT-4 Turbo beats random search 81.25% of the time). We also compute the change in validation error for each optimizer versus random search and report the median and mean change across $32$ tasks. The mean rank is computed between the $5$ HPO approaches and random search, i.e., each method is assigned a rank between $1$ and $6$ on a task. The mean rank for random across the 32 tasks is $4.00$.
@@
Model Versus Random ($\uparrow$) Median Change ($\uparrow$) Mean Change ($\uparrow$) Mean Rank ($\downarrow$)
@@
In Table 1, we report aggregated results comparing two LLMs, random search, and the two BO algorithms. We compare the difference in validation error versus random search ($0.2$ to $0.1$ is a $50\%$ improvement). Every algorithm has a search budget of $10$ evaluations. For random search, we randomly sample $500$ configurations for each model and dataset combination, evaluate the corresponding losses, and then use a bootstrapped sample to estimate the loss on a given budget. As a summary metric, we report the mean and median improvement in the $32$ tasks of each algorithm versus random search. We show the performance on each dataset and model in Figure 7 and Figure 8. GPT-4 Turbo beats random search most frequently, achieving the highest mean and median improvement. When we compare the mean rank for the five search algorithms across the $32$ tasks, GPT-4 Turbo is again consistently better.
Table 2: Chain-of-thought reasoning prompting may offer modest improvement versus prompting a model to immediately output the next set of hyperparameters. We often saw reasoning that describes the effects of adjusting various hyperparameters.
@@
Model Versus Random ($\uparrow$) Median Change ($\uparrow$) Mean Change ($\uparrow$) Rank ($\downarrow$)
@@
In Table 2, we evaluate the effects of chain-of-thought (Wei et al., 2022b) on performance. To elicit reasoning, our intermediate messages to the language model are:
@@
The loss is expressed in scientific notation to four decimal places. See Appendix B for complete details on the prompts. We see that including reasoning has a large positive impact on GPT-3.5 and a marginal impact on GPT-4. Reasoning chains can be informative for a practitioner at the cost of additional tokens. We show a snippet of GPT-4 reasoning when tuning hyperparameters for a neural network:
  • (loss $0.193$) GPT-4: The loss remained the same despite changes in the hyperparameters. This suggests that the model might not be sensitive to these parameters or has reached a local minimum. We should try a more drastic change in the learning rate and depth.
@@
There are potential concerns for the faithfulness of explanations (Lanham et al., 2023), though we observe consistent reasoning that trades off between exploration and selecting a strong final value. We also see logic that understands interactions and effects of different hyperparameters, presumably garnered from the LLM training on textbooks and other resources on HPO. We provide additional full trajectories in Appendix G.
@@
We evaluated our approach on longer time horizons, with up to $100$ proposed configurations, and show the results in Table 3 and Table 5. Here, we used the “compressed” prompt described in Section 2: we prompted the LLM with a single message at each step, showing the entire history of hyperparameter configurations and losses. A longer search trajectory has a reduced dependency on initialization and requires a better search algorithm for good performance. GPT-4 Turbo still achieved a better average rank with hyperparameter optimization trajectories up to length $60$ and is comparable in trajectories of length $100$.
We observe the reasoning generated by the language model is a useful summary of experimental results and the relative importance of different hyperparameters. For example, when tuning a random forest, the LLM outputs the following reasoning:
@@
We additionally ran a preliminary experiment to assess if LLMs can be useful to initialize Bayesian optimization—we find that for a search trajectory of length $30$ total steps, using the GPT-4 Turbo proposed configurations for the first ten steps improves or matches performance on $21$ of the $32$ tasks ($65.6\%$) for Bayesian optimization with a random forest surrogate function. This suggests LLMs can be synergistic with existing approaches.
4.2 CIFAR-10
To evaluate performance in a more difficult deep learning scenario, we tested our method on tuning Vision Transformers (ViTs) (Dosovitskiy et al., 2020) and ResNet-9 (He et al., 2016) on CIFAR-10. This setup poses a challenge for hyperparameter tuning since ViTs typically operate in the regime of large-scale pretraining. We tune five hyperparameters: optimizer choice, learning rate, batch size, weight decay, and label smoothing. To make the task more difficult, we limit training to just 20 epochs and provide minimal context to the LLM, prompting with “You are helping tune hyperparameters for a neural network” without any details about the architecture or dataset. We use the compressed prompting format without chain-of-thought.
Despite these constraints, the LLM successfully optimized validation loss, outperforming random search as shown in Figure 3. Again, we focus on the initial search phase because LLM-based approaches can leverage learned information about hyperparameters to make better proposals. We evaluated two approaches for hyperparameter specification from the LLM: (1) using the same structured search space as random search and (2) allowing free-form valid hyperparameter suggestions, such as requiring batch size to be a positive integer but not restricting its range. The second approach provides more flexibility while maintaining validity constraints. Full experimental details are provided in Appendix C.
Figure 3: Performance comparison of hyperparameter optimization methods on CIFAR-10. Left: Tuning Vision Transformers shows LLM-based approaches achieve lower validation loss compared to random search after 30 iterations. The config-based LLM approach, which uses explicit hyperparameter ranges, performs similarly to the unconstrained LLM. Right: Similar results for ResNet architecture. The best validation loss is tracked across iterations to reflect real-world tuning scenarios.
4.3 Ablation Studies on Prompt Information and Robustness
We conducted experiments on varying the amount of information presented in the prompt while still tuning a Vision Transformer with the hyperparameter search space above. These are plotted in Figure 4.
What is the effect of specifying additional details about the model and dataset?
We investigated how providing varying levels of detail in the initial prompt affects hyperparameter tuning performance for Vision Transformers. In addition to the base prompt, we considered three conditions:
  1. 1.
    Adding dataset information (“…tune hyperparameters for CIFAR-10”)
  2. 2.
    Adding both dataset and correct architecture (“…tune a Vision Transformer on CIFAR-10”)
  3. 3.
    Adding dataset and incorrect architecture (“…tune a convolutional neural network on CIFAR-10”)
While all conditions reach similar performance by iteration 10 (loss $\approx$ 0.65), the amount of information provided can improve hyperparameter selection in the first two steps.
Figure 4: Effect of prompt information and measurement noise on hyperparameter optimization for Vision Transformers. Left: Best validation loss across tuning iterations with varying levels of initial prompt detail, from basic instructions to including dataset and architecture information. The amount of information provided can improve hyperparameter selection in the first two steps but all conditions reach similar performance at iteration 10. Right: Comparison of optimization performance with clean versus noisy (±$\pm 10\%$) loss measurements. The similar performance suggests robustness to measurement noise.
How robust is the LLM to noisy measurements?
We found that LLM performance was similar even when there was noise in the optimization process. Here, we multiplied the loss and accuracy by a random value sampled uniformly from (0.9, 1.1) at each step before passing it to the LLM. This suggests that LLMs can be robust to noisy measurements in the inner optimization loop, which is important since the process of training neural networks can have stochasticity in validation metrics.
Visualizing the Search Trajectories
We randomly sampled two hyperparameter search trajectories and visualized the evolution of the batch size and the learning rate in Figure 5. In the trajectories we examined, we found a reasonable hyperparameter search strategy. There is evidence of conditional hyperparameter selection, where the LLM selects a learning rate for SGD higher than all its choices for Adam, which follows conventional wisdom. SGD is used only once, which may be appropriate in this small-budget problem. Bigger batch sizes are not tried after a batch size of $256$, which resulted in a higher loss.
Figure 5: Visualization of hyperparameter optimization trajectories from the GPT-4 tuning ResNet. GPT-4 selected SGD once (denoted with a plus sign) and Adam with the remaining proposals in both training runs. The learning rate is adjusted accordingly to the optimizer and regions with high loss were not revisited.
@@ 4.4 Code generation
Code generation offers a flexible paradigm to specify the configuration space for hyperparameter tuning. In this setup, we treat the model and optimizer source code as hyperparameters tuned by the LLM. We use code generation to avoid needing to specify a configuration space to search over. The code generation approach is evaluated against random search and LLMs with a fixed, configuration-based search space. The baseline methods are provided with a search space, including the network’s depth and width, batch size, learning rate, and weight decay. To minimize the risk of data leakage issues where the LLM was trained on performant hyperparameter settings for our dataset, we used a Kaggle NYC Taxi dataset (Taxi & Commission, 2023a, b) for our experiments that was released after the knowledge cutoff of the LLM. We provide our prompts and additional details in Appendix D.
In the first tuning iteration, we asked the LLM (GPT-4 Turbo) to output model and optimizer source code, which implicitly requires specifying the hyperparameters. This code is generated as functions that take in hyperparameters as arguments. Hyperparameter tuning is performed in later iterations by asking the LLM to generate function calls with specific arguments. The model is run with the given settings, and the LLM receives the validation loss and the average training losses during each epoch as feedback.
Table 4 reports the minimum test loss achieved using a fixed search budget of $5$ evaluations, simulating the initial search phase results during hyperparameter tuning. We randomly sample $200$ configurations and use a bootstrapped sample to estimate the standard error for a random search. Figure 6 (bottom) reports the minimum test loss found at each step and illustrates that code generation obtains better initial settings than competing approaches.
Figure 6: Test loss trajectory comparison with a standard error over $5$ runs on ResNet-9 on CIFAR-10 (top) and NYC Taxi (bottom). The initial proposal and early iterations from direct code generation and configuration-based LLM search compare favorably to random search.
@@ 5 Discussion
We discuss limitations and explore broader implications and future work.
Possibility of Dataset Contamination.
While some evaluation benchmarks may exist in LLM training data, our prompting process only provides function outputs without revealing benchmark details, reducing the possibility of direct memorization. Previous research suggests that such contamination might not critically influence overall performance (Radford et al., 2021). Building more private data sets unseen by LLM during training would be valuable for future studies or even synthetic datasets like our quadratic experiments.
@@
The exact inference procedures for LLMs like GPT-4 (OpenAI, 2023) are not publicly disclosed, making replicating the results potentially challenging. While setting the sampling temperature to 00 should theoretically yield deterministic outputs, this is not always observed in practice. For future studies, using open source LLMs (Taori et al., 2023; Touvron et al., 2023a, b) to establish a benchmark with reproducible results would be beneficial.
@@
It can potentially be costly to conduct full-scale HPO experiments with GPT-4. For example, executing $32$ tasks with trajectory $10$ using chain-of-thought reasoning on HPOBench costs approximately \$$\$8$ USD as of March 2024. This cost can increase further when using more iterations but can be reduced using the “compressed” prompt. For instance, ten rounds of hyperparameter tuning with the “compressed” prompt and GPT-4, as in the ablation experiments, only costs five cents. These efficiency benefits of using an LLM for HPO could be worthwhile if the cost of the underlying experiment is expensive. We ran some preliminary experiments with the CIFAR-10 setup using Llama3-8B (open-weights) as the LLM and found it could pick initial hyperparameters but was less effective during the search phase.
5.1 Future Directions
Our experiments demonstrate LLMs’ quantitative and qualitative effectiveness for HPO but also pose a new conundrum for how this ability emerges. Wei et al. (2022a) suggest one hypothesis for emergent abilities: “more parameters and more training enable better memorization, which could be helpful for tasks requiring world knowledge.” This may partially explain the jump in performance we observe from GPT-3.5 to GPT-4, but the underlying mechanisms for this emergence are not yet well understood.
LLMs as Research Assistants
Initial experiments demonstrated promising applications beyond simple hyperparameter suggestion. While traditional optimization methods focus solely on parameter selection, as shown in Figure 10, LLMs provided useful feedback for error messages through natural language interaction. However, models can face characteristic challenges of current language models, including hallucinations (Bowman, 2023). For instance, in Figure 11, we observed the LLM making incorrect assumptions about model architecture and suggesting potentially inappropriate regularization strategies.
Our findings suggest several promising research directions. First, exploring how to ground LLM suggestions in verified technical knowledge could improve reliability. Second, developing more robust dialogue systems for experimental debugging could enhance the assistant paradigm. The future may lie in more interactive and intuitive optimization processes that maintain human oversight while making use of the general knowledge of language models.
5.2 Ethics Statement
The approach presented in this paper facilitates machine learning research and applications by potentially making it easier to find well-performing hyperparameters. Using LLMs as hyperparameter tuners may help lower the cost of model training (e.g., time, compute, and environmental impact), making machine learning experiments easier for those from other disciplines. Overall, the benefits and risks are likely similar to those of other automatic machine learning (AutoML) research. Note also that there are potential risks to code generation with LLMs; Zelikman et al. (2023) present a recent overview.